Circular No. 6837

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1997 XF_{11}

This object, discovered by J. V. Scotti in the course of the Spacewatch program at the University of Arizona on 1997 Dec. 6 (cf. MPEC 1997-Y11) and with H = 17.0 recognized as one of the 108 “potentially hazardous asteroids” (PHAs), has been under observation through 1998 Mar. 4, the latest data having been obtained by P. J. Shelus with the 0.76-m relector at the McDonald Observatory. An orbit computation from the 88-day arc appears on MPEC 1998-E13 and MPC 31424. This nominal orbit indicates that the object will pass only 0.00031 AU from the earth on 2028 Oct. 26.73 UT! Error estimates suggest that passage within 0.002 AU is virtually certain, this figure being decidedly smaller than has been reliably predicted for generally fainter PHAs in the foreseeable future. The following ephemeris is given in the hope that further observations will allow refinement of the 2028 miss distance; physical observations would also presumably be useful. The object will next be at opposition in Feb. 2000 at V = 19.3; there will also be an approach to a distance of 0.065 AU on 2002 Oct. 31 (V = 13.7).

1998 TT α_{2000} δ_{2000} Δ r ϵ β V
Mar. 8 6^{h}15.^{m}92 +17^{o}56.^{′}7 1.557 2.070 106.^{o}5 27.^{o}4 20.8
18 6 22.81 +18 18.0 1.701 2.089  98.2 28.1 21.0
28 6 32.08 +18 33.4 1.846 2.106  90.5 28.3 21.2
Apr. 7 6 43.26 +18 42.1 1.988 2.119  83.3 28.0 21.4
17 6 55.94 +18 43.5 2.126 2.128  76.5 27.3 21.5
27 7 09.81 +18 36.9 2.258 2.135  70.0 26.3 21.6
May 7 7 24.64 +18 22.1 2.382 2.139  63.9 25.1 21.7
17 7 40.21 +17 58.8 2.497 2.139  57.9 23.6 21.8
27 7 56.37 +17 26.9 2.602 2.136  52.2 22.0 21.8
June 6 8 13.00 +16 46.4 2.695 2.131  46.7 20.3 21.8
16 8 30.00 +15 57.3 2.777 2.122  41.3 18.4 21.8
26 8 47.30 +14 59.7 2.845 2.110  36.0 16.5 21.8
July 6 9 04.85 +13 54.0 2.901 2.094  30.9 14.4 21.7

COMETS C/1997 W2 AND C/1998 E1 The SOHO-LASCO Consortium has reported two more sungrazing

comets. Further details are given on MPEC 1998-E21 and 1998-E22.

UT  α_{2000}  δ_{2000}
1997 Nov. 19.360  15^{h}33.^{m}4  −21^{o}54′  C/1997 W2
1998 Mar. 2.985  23 12.8  − 7 44  C/1998 E1
1998 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden